A chemist has a solution that is 60% saline and another solution that is 35% saline. She wants to make 100 ounces of a mixture that will be 40% saline. How many ounces of the 60% saline solution will she use?
Answer by mananth(16946)

percent ---------------- quantity
Saline solution I 60.00% ---------------- x ounces
Saline solution II 35.00% ------ 100 - x ounces
Mixture 40.00% ---------------- 100
12
60.00%x +35.00% (100 -x)=40.00%*100
60x+35 (100-x) =4000
60x+3500-35x = 4000
60x- 35x = 4000 -3500
25 x = 500
/ 25
x=20 ounces 60.00% Saline solution I
80 ounces 35.00% Saline solution II

m.ananth@hotmail.ca